Find the balance of the account after;
2 years if K3000 is invested at 8% per annum.interest compounded quarterly
b)5years if K2000 is invested at 6% per annum,interest compounded 6 monthly.
c)4years if K5000 is invested at 12% per annum,interest compounded monthly.
d)5 1/2years if K960 is borrowed at 11% per annum,interest compounded fortnightly.

Answer by Theo(13342)

Find the balance of the account after:

a) 2 years if K3000 is invested at 8% per annum compounded quarterly

2 years * 4 quarters per year = 8 quarters.
8% per year / 4 quarters per year = 2% per quarter.
3000 * 1.02 ^ 8 = 3514.978143.

b) 5 years if K2000 is invested at 6% per annum compounded 6 monthly.

5 years * 2 semi-annual time periods per year = 10 semi-annual time periods.
6% per year / 2 semi-annual time periods per year = 3% per semi-annual time period.
2000 * 1.03 ^ 10 = 2687.832759.

c) 4 years if K5000 is invested at 12% per annum compounded monthly.

4 years * 12 months per year = 48 months.
12% per year / 12 months per year = 1% per month.
5000 * 1.01 ^ 48 = 8061.130388.

d) 5 1/2 years if K960 is borrowed at 11% per annum compounded fortnightly.

5.5 years * 26 fortnights per year = 5.5 * 26 = 143 fortnights.
11% per year / 26 fortnights per year = .4230769231% per fortnight.
960 * 1.004230769231 ^ 143 = 1755.759973.

let me know if you have any questions.
theo